OpenStudy (anonymous):

hi, how do I answer a question where i (imaginary number) has a negative exponent?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

example?

OpenStudy (anonymous):

i^(-1)=-i take the minus out of the exponent and attach it to i.

OpenStudy (anonymous):

\[i^{-1}=\frac{1}{i}=\frac{1}{i}\times \frac{i}{i}=-i\]

OpenStudy (anonymous):

but i do like saubhik's explanation!

OpenStudy (anonymous):

thank you guys sooo much :D

OpenStudy (anonymous):

use the "good answer" button:)

OpenStudy (anonymous):

i do not know where that is :/

OpenStudy (anonymous):

see file..

OpenStudy (anonymous):

thank you but it doesn't appear on my screen. should i log out and than log back in do you think?

OpenStudy (anonymous):

yeps..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you both again :)